OpenStudy (anonymous):

Vertical asymptotes of the rational function for: f(x)=x+9 divided by x^2-64 thanks, a medal awaits!

OpenStudy (anonymous):

Youmean this function? \[f(x)=\frac{ x+9 }{ x^2-64 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First take a look at the points where f(x) is not defined Can you tell me that point

OpenStudy (anonymous):

you must find a point where f(x) is losing continuity

OpenStudy (anonymous):

eh...0

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

replace x with 0 and see what f(x) is

OpenStudy (anonymous):

-8,8

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

oh I see what you mean because you can factor the denominator right?

OpenStudy (anonymous):

then you have a domain to settle

OpenStudy (anonymous):

still learning my precalc :)

OpenStudy (anonymous):

the numbers must be greater than 0

OpenStudy (mathmale):

Please just factor the denominator of \[f(x)=\frac{ x+9 }{ x^2-64 }\]and then let each of the factors of the denominator equal zero. Write the results as equations: x = a, x = b, and they will be the equations of your vertical asymptotes.

OpenStudy (anonymous):

like (x-8)=0 and (x+8)=0

OpenStudy (mathmale):

@Opulentgal : The best way to re-state your assertion, "the numbers must be greater than 0," is to state: "The denominator cannot be zero / is undefined if the den. is zero." @Opulentgal : that's great! Now please solve each of those equations for x. Remember, you must write the results as equations.